Fermi National Laboratory


Accelerator Update

Wednesday May 19
The day shift began with Operations monitoring store 3510, with the Antiproton Source (Pbar) stacking, with Recycler experts conducting studies, and with MiniBooNE taking beam.

At 8:44 AM, the Collider Detector at Fermilab (CDF) called asking that the Tevatron (TeV) beam gap problem get fixed. Operators and TeV experts tuned and adjusted collimators. CDF was able to turn back on by 9:47 AM.

Operations established beam to the Meson MTest experiment at 9:24 AM.

Pbar experts began a stacktail study at 9:54 AM, followed by a Pbar Debuncher RF study about an hour later.

Booster experts began an extraction study at 10:01 AM.

A Main Injector (MI) expert began a study period at 12:27 PM. This study interfered with both Switchyard (SY) and MiniBooNE beam.

At 2:10 PM, a Linac RF station (LRF2) tripped off for the first of four times in eight minutes. It eventually reset okay.

An MI expert called at 2:41 PM and said that in order to resolve a MI-8 vacuum problem they would have to close a beam valve for ten minutes. The MI-8 beamline transports beam from the Booster to MI.

At 6:46 PM, a Pbar Debuncher cryo system wet engine tripped off. Central Helium Liquefier (CHL) operators called in cryo system electrical technicians to investigate.

At 7:40 PM, a TeV power supply (T:VE16) tripped off and caused store 3510 to abort. There was no quench. Operators replaced the regulator for this power supply.

Operators prepared the D-zero (D0) experiment for a one-hour collision hall access at 8:19 PM.

The MTest experiment finished using beam for the night at 8:19 PM.

Cryo system techs completed their work on the Pbar wet engine at 9:55 PM, but a problem with its regulation control loop meant it had to be run in local at a specific speed. Experts will continue their work in the morning.

Operations put the TeV in shot setup at 10:02 PM.

Operations established store 3513 at 11:56 PM with an initial luminosity of 61.9E30.

Thursday May 20
The midnight shift began with Operations monitoring store 3513, with Pbar stacking, and with MiniBooNE taking beam.

Operations prepared MCenter for beam at 12:21 AM.

A Recycler expert began stashing antiprotons from Pbar at 4 AM.

At 5:49 AM, some TeV sextupoles power supplies tripped off and caused the store to abort. There was no quench.

Operation put the TeV through a dry squeeze at 6:10 AM.

MCenter finished using beam at 6:26 AM. Operations turned off all SY beam.

TeV experts began a study period at 7:42 AM.

Operations put the TeV through a dry squeeze at 9 AM.

Cryo system experts fixed the Pbar wet engine controls problem at 10:42 AM.

Operations established store 3516 at 11:37 AM with an initial luminosity of 50.5E30.

Operations had Booster ready for access by 11:37 AM.

At 12:51 PM, the Booster Debuncher charging supply tripped off. Operators had the Debuncher back on within twenty-five minutes.

Operations reported at 12:58 PM that the Booster chopper was tripping off on a power supply interlock indication. A Booster expert investigated, but couldn't find anything wrong. It tripped off twice more, but then settled down and was running normally by 2:31 PM.

Pbar stacking resumed at 1:23 PM.

Operations established beam to MTest at 1:50 PM.

At 2:05 PM, a Booster power supply (B:CEPSE1) began tripping off repeatedly. Experts investigated, but could find nothing wrong. It stopped tripping. At 3:34 PM, Booster experts began preparing a new supply.

Pbar resumed stacking at 2:45 PM.

An MI expert began a regulation study at 3:05 PM.

A TeV expert reported at 3:35 PM that the TeV beam gap problems were due to a problem with the longitudinal dampers.

MTest finished using beam for the night at 6 PM.

The Booster power supply CEPSE1 began tripping off again at 6:11 PM. Operators replaced the power supply.

MI experts began a study period at 8:52 PM.

Friday May 21
The midnight shift began with Operations monitoring store 3516, with Pbar stacking, and with MiniBooNE taking beam.

At 1:14 AM, TeV experts began a ten-minute TeV RF (TRF7) study. CDF and D0 turned off their detectors for this study.

Operations established beam to MCenter at 1:15 AM.

Operations terminated store 3516 at 4:30 AM and halted stacking for shot setup.

Operations put the TeV into shot setup at 5 AM.

At 6:58 AM, shortly after the last antiprotons were loaded into the TeV for the store, the TeV suffered a sector A1 and F4 quench. An expert began his investigation. Pbar resumed stacking.

Plans for Today
The plans for today include allowing CDF and D0 to access their collision halls this morning. TeV experts will conduct studies during the day until the stack reaches a level sufficient enough to establish a store. The plan for this weekend is to stack and store.

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.



last modified 5/21/2004   email Fermilab